# Visual Affordances: Enabling Robots to Understand Object Functionality

### v1

@@ II-D Functional classification
One of the main challenges of affordance classification is that the same object in the scene has multiple affordances that vary depending on the task. For example, a cup on a table can suggest the action of picking or filling, but until a task is defined (e.g. ‘move the cup’), both affordances are plausible. Another challenge is the fact that objects with similar appearances might afford different actions. For example, some models of trowel and turners might be similar in colour and shape, however the surface of a trowel is used to scoop, while the surface of a turner to support.
TABLE II: Characteristics and comparison of methods for function classification (or also known as affordance classification). Note that these methods use either object detection, object segmentation, object classification, or their combination as auxiliary tasks.
@@
We summarise the characteristics of affordance classification methods in Table II. Methods learn actions that can be performed with objects in the scene either from human demonstration [3, 58], or from images of the environment [52, 54]. Sun et al. [52] used Probabilistic Graphical models to relate affordances with the appearance of objects encoding the image using Principal Component Analysis. The main limitation of the method is the limited scalability to high-dimensional signals like images and the increasing complexity of the graph structure when adding affordance categories. To limit the input dimensionality and focus only on regions of interest in the image, Zheng et al. [54] use object detection before predicting the action related to object regions. Moreover, the combination of affordance classification with auxiliary tasks such as detection and segmentation allows to group objects based on the actions they are used for (functionality), instead of their appearance [3, 54]. By training methods on data of people using objects or with the robot exploring the environment, previous works [3, 52, 58] implicitly considered as a task the functional use of the object. These methods do not enable the physical interaction of a robotic hand with the object, as the action is not associated with an interaction region in the image [52, 54, 58]. This results in the robot having multiple options (ambiguity) on how and where to perform the interaction. For example, even a simple instruction like “move the cup” can suggest a robot multiple ways to perform the interaction, such as grasping the cup by the body or by the rim.
TABLE III: Characteristics and comparison of visual affordance segmentation models [4]. Note that we report the best-performing backbone for each model, and we do not consider additional parts of the pipelines, such as a separate object detector.
Model Architecture Attention Aff. Object CRF
Source Backbone FPN IF Sp Ch Sa Mc C E C S L
@@
KEYS – Source: reference architecture, Backbone: visual encoder, Sp: spatial attention, Ch: channel attention, Mc: masked cross-attention, Sa: self-attention, Aff.:  affordance, C: classification, E: edge segmentation, S: segmentation, L: localisation, RN: ResNet, RNX: ResNeXt, RNS: ResNeSt, SE-RNX: squeeze and excite ResNeXt, DRN: Dilated Residual Network, CRF: conditioned random fields, IF: intermediate feature maps fusion; $\bullet$: considered, $\circ$: not considered.
II-E Functional segmentation
The segmentation of functional regions on objects in the image identifies where the robot needs to perform the interaction with the object and is approached in the literature in two main ways (see Fig. 4): affordance detection and segmentation aims at finding the objects of interest in the image and separating the functional regions on the object, and affordance grounding details the task to accomplish and identifies the region on the object that should be used to perform the action.
Affordance detection and segmentation. Given an image $I$, the model predicts bounding boxes $\{b_{o}\}^{O}_{o=1}$ and segmentation masks of $A$ functional regions $\{S_{o}\}^{O}_{o=1}$ for a set of objects of interest,
@@ II-E Functional segmentation
The segmentation mask $S_{o}$ can be also formulated as the combination of the actions $\{c_{a}\}^{A}_{a=1}$ with a probability map $\{S_{o,a}\}$ where $S\in[0,1]^{W\times H}$ indicates the region where an action takes place for each object [20, 86]. Affordance detection and segmentation formulation assumes that the objects of interest are the ones annotated in the dataset, that the task $\mathcal{T}$ is to use the object to fulfil the purpose it was designed for  [9, 10, 12, 62], and that different parts of the objects are associated with a functionality to accomplish the task. For example, a knife is designed to cut another object and therefore the handle is designed to be grasped while the blade is used for the cutting. Therefore, the models should localise the object that offers “cut” as an affordance (affordance detection) and segment the blade as a part of the knife offering the action “cut” (affordance segmentation).
@@
Previous methods [63, 65, 79, 67, 66, 64, 61] adapt semantic and instance segmentation architectures to predict affordance regions on the objects. For example, A-Mask R-CNN [73] and AffordanceNet [10] modify an instance segmentation model (Mask R-CNN [67]) to predict the affordance masks instead of the object masks for each object localised in an image. Starting from the design of AffordanceNet, BPN [83] and 4C-RPN-5C [68] align the region of interest with the feature maps at different resolutions and predict the overlapping of bounding boxes and boundaries of affordance regions. An off-the-shelf object detector localises regions of interest in the image, however inaccurate or wrong predictions can consequently result in segmenting affordance regions within image parts outside of the actual objects. When edges are blurred or not clearly defined (e.g. occlusions or transparent objects), BPN fails to predict precise affordance contours despite its additional affordance edge segmentation component in the model architecture. Instead, semantic segmentation models [17, 20, 62, 78, 81, 84, 74] avoid the dependence from an object detector and assign each pixel of the image to an affordance class during supervised training (per-pixel affordance segmentation). When objects are occluded or boundaries are not clearly defined, methods such as CNN [62], RN50-F [17], and ACANet [20] can classify affordance pixels outside the object region.
@@
Attention mechanisms [83, 78, 81, 84, 74] are an alternative way to consider only relevant information in the image by weighing feature maps extracted from the image. For example, GSE [74] learns the channels weight with supervision of the affordances. DRNAtt [78], SEANet [81], and BPN [83], learn similarities between channels or positions in the feature maps without direct supervision, updating the weights of the layers that compute the attention map during training. For computational reasons, both DRNAtt and GSE process feature maps at low-resolutions where important details (e.g. edges) for affordance segmentation are degraded when the object is not close to the camera. In RANet [84], the attention weights are learned with the supervision of object classes. However, in case of occlusions, mistakes in the attention weights cause the mismatch between the predicted object classes and the segmented affordances.
@@
Previous methods [17, 20, 62, 78, 81, 84, 74] coupled the classification of the affordances and the segmentation of regions in the image. However, the two subtasks can be decoupled by performing class assignment at the level of each segmentation mask [86]. For example, STRAP [86] is designed as a multi-branch architecture that learns the affordance classification in one branch and the segmentation masks in another branch. The model is train to segment the affordance mask with a weakly supervision from a point annotation of a region [89] and by using Conditional Random Fields that process the pixel position and colour. However, this approach can lead to inaccurate segmentations when the object colour is not clearly distinguished from the background [16, 62]. STRAP also uses self-attention to process low-resolution feature maps extracted by the backbone, losing details about the object in the image when the object scale is small. To increase the resolution of processed feature maps, M2F-AFF [4] adapts Mask2Former [88]. Mask2Former added masking in the cross-attention mechanism to combine the features extracted from the image with learnable latent vectors, while ignoring the pixel positions outside the object region (background).
@@
Table III summarises the characteristics of the methods we discussed for affordance detection and segmentation. Note that the segmentation of affordance regions is tackled independently from the end-effector. Affordance regions can be mapped to the robot control process to perform the actions [10, 12, 83], yet they represent a coarse information, since characteristics of end-effectors such as the number of fingers or the degrees of freedom influence the contact regions on the object (more fine-grained than affordance regions).
@@
Affordance grounding. Given an image $I$ and a task $\mathcal{T}$, the model predicts the probability map $\{S_{o}\}^{O}_{o=1}$ to identify the image region that the agent uses to interact with the object,
@@
The task $\mathcal{T}$ can be expressed through natural language [90], an affordance category [11, 91, 92, 93], a point in 2D [94, 95], or another image of the object of interest [96, 97, 98]. With this formulation, one-shot methods that use prior information (e.g. an image) can also be seen as affordance grounding methods.
TABLE IV: Characteristics and comparison of affordance grounding methods.
@@
We summarise the characteristics of affordance grounding methods in Table IV. Most of the methods use the action to be performed on the object as a prior [11, 91, 92, 95], however the prior for the affordance grounding can be a 2D query point in the image plane [94], a query image [96, 97, 98], or the task in natural language [90]. These priors enable affordance grounding methods to tackle generalization to different object categories while avoiding an explicit object detection phase. One approach to perform affordance segmentation is through weak supervision i.e., using methods for explainability for example Class Activation Maps (CAM) [99] to obtain the region in the image that corresponds to the action [11, 91, 96]. The use of CAM, which highlights coarse image regions not bounded by object contours, limits the application of these methods to unoccluded object settings. One-shot-based methods use an image as a prior to select objects of interest [19, 100], or segment affordance regions [97], based on the similarity between the input images and the prior (query image). However, the support image is assumed to be similar to the query images, thus implying that the object category in the scene should be known in advance.
@@
To cope with the limited amount of training images, methods learn affordances using pretrained multimodal models [90, 92, 93, 95], using knowledge transfer from video to images [91, 98] or from exocentric views of objects to the egocentric ones [11, 96]. In particular, multimodal models help generalising to unknown object categories or unknown actions (open vocabulary). For example, AffordanceCLIP adapts CLIP [44] with a learnable feature pyramid network to predict the probability map [95]. A contrastive loss encourages pixel-level embeddings within the annotated mask of the object to align with the corresponding linguistic feature. AffordanceLLM [90] processes vision and language information using a Large Language Model (LLM) to predict tokens for the affordance segmentation. This LLM generates text tokens that are converted into a description of the part used to perform the task and a mask token that is combined with the visual tokens using a transformer decoder to predict the affordance map. Few of the methods [11, 96] for affordance grounding focused on learning object affordances building correspondences from the exocentric view of an object (human using the object) to the egocentric one (object only). Both LOCATE [96] and Cross-View-AG [11] during training combine a loss to learn the affordance category with strong supervision with losses to preserve the similarity between the outputs of the exocentric and egocentric images processing. Instead of learning directly from images, methods like Demo2Vec [98] and Hotspots [91] learn to transfer the affordance from videos of humans interacting with objects in household environments e.g., oven, fridge, washing machine, to the images containing only the objects. Although these methods [11, 91, 96, 98] can learn the object affordances from the information of how humans performs actions with the object, the egocentric views of the objects are composed by the object on a plain background without occlusions or clutter, limiting the generalisation to in-the-wild images.
@@
Despite generalising to different actions or task formulations, affordance grounding methods output confidence maps that are not bounded by object edges. The confidence maps could also overlap with other objects in case of clutter or with a human hand if the object is hand-held. Using a coarse confidence map when interacting with an object can lead a robot to misplace the end-effector, thus undermining the success of the interaction or harming the human.
II-F End-effector pose estimation and synthesis
@@
To perceive the visual affordance, the robot needs to predict also how the interaction with the object will be performed i.e., the pose of the end-effector on the object. Previous works [2, 14, 23, 24, 25, 101, 102, 103, 104] related the problem mostly to grasping rather than to visual affordances, and referred to the problem differently depending on the characteristics of the end-effector: grasping detection in case of a two-fingers gripper [2, 23, 24, 25, 103], hand-object pose estimation in case of the human hand [5, 14] (e.g. MANO model [31]), multi-finger grasping in case of the three fingers Barrett hand [101]. Since also the hand can be considered as a multi-finger end-effector, we merge the discussion on three fingers and human hand into multi-finger pose estimation and interaction synthesis. Once the end-effector model is chosen, and given the image of the object and the task, the model predicts the pose $P$ of the end-effector on the object,
@@
Predicting the pose of the end-effector, however, is challenging because the end-effector is not observed in the image, and therefore only the visual features of the object can be used.
Figure 5: The pose of a two-fingers gripper encoded as an oriented rectangle on the image plane [24]. The 5 dimensions of the rectangle encode: the centre ($x$, $y$), the opening ($w$), the fingers length ($h$), and the orientation $\theta$.
Grasping detection. Assuming a two-fingers gripper as end-effector, the pose estimation of the end-effector on the object can be reformulated as a detection of grasping points directly on the image, encoding the parameters of the gripper as a rectangle [2]. The poses of the gripper on the object are detected using an oriented rectangle representing the 5 DoFs of a parallel plate gripper on the image plane (see Fig. 5): 1 DoF encodes the gripper rotation with respect to the horizontal axis, 2 DoFs encodes the translation of the gripper centre (on the horizontal and vertical axis), and 2 DoFs encodes the geometry of the gripper (1 for the opening width and 1 for the fingers length). The underlying assumption is the availability of a depth map to obtain the full 7 DoFs representation of the gripper in 3D (3 DoFs for translation, 3 DoFs for rotation, and 1 DoF for the opening width). Given an RGB-D image $I\in\mathbb{R}^{W\times H\times 4}$, the model predicts a set of $G$ oriented rectangles $\{r_{g}\}^{G}_{g=1}$ with $r\in\mathbb{R}^{5}$ consisting of the coordinated of the rectangle centre, the rectangle size (width and height), and the orientation. Predicting the pose of a two-fingers gripper on the surface of an object is challenging because each object has multiple grasping points, but only a part of grasping poses leads to a successful grasping. Moreover, when estimating the grasping points from a single view, only a side of the object is visible, limiting the number of feasible grasping points. Such a formulation, i.e. predicting an oriented rectangle, corresponds to the formulation of object detection and therefore object detection models, pre-trained on large-scale datasets, can be re-used and retrained or fine-tuned for grasping detection [24, 104]. However, the oriented rectangle representation can lead to a wrong detection of the grasping points if the object is in a challenging pose, and consequently the robots can fail to grasp the object.
TABLE V: Characteristics and comparison of grasping detection models.
@@
Table V summarises methods for grasping detection. Most of the methods [2, 23, 24, 25, 103] use RGB-D images to predict grasping rectangles, as the depth information provides geometric cues to the model. Different ways of fusing visual information include learning to fuse the information in the first layers of the method  [2, 23, 24, 103] (early fusion), or using a separate backbone to process RGB and depth before fusion (middle fusion) [25]. However, there are no results showing that a fusion mechanism is more effective than the others. The feature extraction is performed mainly convolutional networks like ResNet [23, 24, 25, 104] or AlexNet [102] pre-trained on ImageNet [105], to transfer the feature representations learned on large scale datasets. Similarly to object detection, methods can be categorised into single-stage and and two-stage approaches. Single-stage methods [25, 102, 103] predict the final oriented rectangles from the image, either directly regressing the rectangle [25, 102] or considering the rectangle as a by-product of object segmentation [103]. Two-stage methods [2, 23, 24, 104] first predict grasping candidates (coarse estimation) and then refine the predictions (fine estimation). The majority of two-stage methods adapt works for object detection (e.g. Faster R-CNN [55]) to grasping detection in different ways: separating the learning of the object location and grasp locations [24]; separating the learning of the quantized angle from the learning of the centre, width and height of the grasping rectangle [23]; separating the coarse prediction of grasping rectangles from the refinement based on the semantic segmentation of the object [104]. Auxiliary tasks, such as object detection [23] and segmentation [104], can help constrain the prediction of the grasping rectangle to the part of the image containing the object, reducing mistakes in case of cluttered scenes or when the object is not in foreground and completely visible. Other auxiliary tasks are the likelihood of an image patch (non-overlapping piece of the image) containing a grasp [102] limiting the predicted grasping rectangles number to some parts of the image, and the grasping region segmentation [103] constraining the grasping rectangle to graspable region of the object e.g., the handle of a spoon.
@@
Such a formulation of grasping detection lacks a definition of the visual affordance, resulting in non-physically plausible solutions as the robot can grasp the object at any location on the object’s surface. For example, the rim of a cup filled with liquid (suggesting the affordance of pouring the content) might be selected as a potential grasping point without considering that the liquid might be spilled or damage the end-effector. Most of the methods for grasping detection [2, 23, 24, 25, 102, 106] use the assumption that objects are observed on a tabletop or on the floor (top-down camera view). Hence, models may fail to generalise to scenarios with different camera view-points or with occlusions (e.g. hand-occlusion in human-robot collaborations).
@@
Multi-finger pose estimation and interaction synthesis. Previous works [5, 14, 101] considered as visual affordance the pose of an end-effector, with more than two fingers, on objects in the scene. Given an image $I$, the model predicts the 6D pose of the end-effector on the object $\{[R|T]_{o}\}^{O}_{o=1}$ with $[R|T]$ representing pose of the end-effector (pose estimation) and renders an image of the end-effector, $\tilde{I}\in\mathbb{R}^{W\times H\times 3}$ (interaction synthesis).
@@
TABLE VI: Characteristics and comparison of multi-finger pose estimation and interaction synthesis methods.
@@
Table VI compares the characteristics of methods for multi-finger pose estimation and interaction synthesis. Methods are based on a coarse-to-fine approach that first locates where the end-effector is supposed to interact with the object and then refines the pose using generative adversarial networks [14, 101] or diffusion models [5]. GanHand [14] estimates objects’ shapes and locations using a sub-network, either an object 6D pose estimator or a reconstruction network. The predicted shape is projected onto the image plane to obtain a segmentation mask that is concatenated with the input image and fed to a second sub-network that predicts the grasp type, i.e. the type of interaction between end-effector and object. A module predicts the coarse pose of the end-effector from the grasp type and visual features. The network refines the end-effector parameters and obtains the final shapes and poses (i.e. MANO model [107]) by minimising an adversarial loss with a discriminator. Multi-FinGAN [101] adapts GanHand architecture to perform the pose estimation of the Barrett end-effector on the object in the image. In Multi-FinGAN the reconstruction of the object is not projected on the image to focus only on the object region, but only to refine the coarse pose of the end-effector. As a consequence, the method underperforms if multiple objects are present in the scene. AffordanceDiffusion [5] uses two diffusion models to generate the image of the end-effector interacting with the object in the image. The diffusion process is based on a end-effector prior (forearm mask) that is composed by a circle representing the end-effector and a rectangle representing the forearm. For every diffusion step, the first model (i.e. LayoutNet [5]) predicts the denoised forearm mask from the features of the forearm mask obtained in the previous step, the object image, and the forearm mask projected on the object image. The predicted layout mask (prior) and the object image are used as input to the second model (i.e. ContentNet [5]) that synthesises an image of the interaction between end-effector and object.

### v2

@@ III-B Functional classification
One of the main challenges of affordance classification is that without a defined task, one object has multiple affordances. For example, a cup on a table can suggest the action of picking or filling, but until a task is defined (e.g. ‘move the cup’), both affordances are plausible. Another challenge is that objects with similar appearances might afford different actions. For example, some models of trowel and turner might be similar in colour and shape, however the surface of a trowel is used to scoop, while the surface of a turner to support.
TABLE II: Comparison of methods for functional classification. Note that these methods use auxiliary tasks such as object detection or classification.
@@
TABLE III: Comparison of visual affordance segmentation models [4]. We report the best-performing backbone for each model and do not consider additional parts of the pipelines, such as a separate object detector.
Model Architecture Attention Affordance Object CRF
Source Backbone FPN IF Sp Ch Sa Mc CLA ES CLA SEG LOC
@@
KEYS – Source: reference architecture, Backbone: visual encoder, Sp: spatial attention, Ch: channel attention, Mc: masked cross-attention, Sa: self-attention, CLA: classification, ES: edge segmentation, SEG: segmentation, LOC: localisation, RN: ResNet, RNX: ResNeXt, RNS: ResNeSt, SE-RNX: squeeze and excite ResNeXt, DRN: Dilated Residual Network, CRF: conditioned random fields, IF: intermediate feature maps fusion; $\bullet$: considered, $\circ$: not considered.
Methods for affordance classification learn actions that can be performed with objects in the scene either from human demonstration [3, 43, 50], or from images of the environment [44, 46]. We summarise the characteristics of these methods in Table II. Nagarajan et al. [43] trained an affordance classifier to predict all the potential actions that a person can perform in an environment (e.g., a kitchen sink). Sun et al. [44] used Probabilistic Graphical models to relate object affordances with appearance. Images are processed with dimensionality reduction, limiting the scalability of the method to high resolution images, and increasing the complexity of the graph structure adding affordance categories. The combination of affordance classification with auxiliary tasks such as detection and segmentation allows to focus only on regions of interest in the image and to group objects based on the actions they are used for (functionality), instead of their appearance [3, 46]. By training methods on data of people using objects or with the agent exploring the environment, previous works [3, 43, 44, 50] implicitly considered as a task the functional use of the object. However, these methods do not consider the physical interaction between the agent and the object, as the action is not associated with an interaction region in the image [44, 46, 50]. This results in the agent having multiple options (ambiguity) on how and where to perform the interaction. For example, even a simple instruction like “move the cup” can be performed in multiple ways, such as grasping the cup by the body or by the rim.
III-C Functional segmentation
The segmentation of functional regions on objects in the image identifies where the agent needs to perform the interaction with the object. This is approached in two ways (see Table III). Affordance detection and segmentation detects the objects of interest in the image and separates the functional regions. Affordance grounding identifies on the object the region that should be used to perform the action defined in the task.
Affordance detection and segmentation. Given an image $I$, the model predicts bounding boxes $\{b_{o}\}^{O}_{o=1}$ and segmentation masks of $A$ functional regions $\{S_{o}\}^{O}_{o=1}$ for objects of interest,
@@ III-C Functional segmentation
The segmentation mask $S_{o}$ can be also formulated as the combination of the actions $\{c_{a}\}^{A}_{a=1}$ with a probability map $\{S_{o,a}\}$ where $S\in[0,1]^{W\times H}$ indicates the region where an action takes place for each object [59, 80]. Affordance detection and segmentation methods assume that the objects of interest are the ones annotated in the dataset, that the task $\mathcal{T}$ is to use the object to fulfil the purpose it was designed for  [6, 11, 12, 54], and that different parts of the objects are associated with a functionality to accomplish the task. For example, in a knife the handle is designed to be grasped while the blade is used for cutting another object. These methods localise the object that affords “cut” (detection) and segment the blade that affords the action “cut” (segmentation).
@@
Previous methods [53, 55, 57, 58, 60, 61, 73] adapt semantic and instance segmentation architectures to predict affordance regions on the objects. For example, A-Mask R-CNN [67] and AffordanceNet [12] modify an instance segmentation model (Mask R-CNN [61]) to predict affordance masks instead of object masks for each localised object. Starting from the design of AffordanceNet, BPN [77], and 4C-RPN-5C [62] combine the region of interest with the feature maps at different resolutions and predict the overlapping of bounding boxes and boundaries of affordance regions. The object detection branch localises regions of interest in the image, but inaccurate or wrong predictions can consequently result in segmenting affordance regions outside of the actual objects. When edges are blurred or not clearly defined (e.g. occlusions or transparent objects), BPN fails to predict precise affordance contours despite its additional edge segmentation component. On the contrary, semantic segmentation models [54, 56, 59, 68, 72, 75, 78] avoid the dependence from an object detector and assign each pixel of the image to an affordance class (per-pixel affordance segmentation). When objects are occluded or boundaries are not clearly defined, methods such as CNN [54], RN50-F [56], and ACANet [59] can classify affordance pixels outside the object region.
@@
Attention mechanisms [68, 72, 75, 77, 78] are an alternative way to consider only relevant information in the image by weighing image feature maps. For example, GSE [68], DRNAtt [72], SEANet [75], and BPN [77], learn the channels weight or the similarity between positions in the feature map without direct supervision. For computational reasons, both DRNAtt and GSE process feature maps at low-resolutions where important details for affordance segmentation (e.g. edges) are degraded for objects not in foreground. In RANet [78], the attention weights are learned with the supervision of object classes. However, in case of occlusions, mistakes in the attention weights cause mismatch between the predicted object classes and the segmented affordances.
@@
Most of previous methods [54, 56, 59, 72, 75, 78, 68] performed the classification of the affordances and the segmentation of regions jointly. However, the two subtasks can be decoupled assigning an affordance class to each segmentation mask [80]. For example, STRAP [80] learns the affordance classification and segmentation in separate branches. The model learns to segment affordance masks with weak supervision from a point annotation of each region [83] and by using Conditional Random Fields to process the pixel position and colour. However, this approach can lead to inaccurate segmentations when the object colour is not clearly distinguished from the background [10, 54]. STRAP also uses self-attention to process low-resolution image feature maps, losing details about the object in the image when the object scale is small. To increase the resolution of processed feature maps, M2F-AFF [4] adapted Mask2Former [82] that combines the image features with learnable latent vectors, while ignoring the pixel positions outside the object region (background) through masked cross-attention.
@@
Note that affordance segmentation is tackled independently from the agent’s hand characteristics, even if the number of fingers or the degrees of freedom influence the contact regions on the object. Nevertheless, affordance regions can be used by an agent such as a robot to perform actions [6, 12, 77].
@@
Affordance grounding. Given an image $I$ and a task $\mathcal{T}$, the model predicts the probability map $\{S_{o}\}^{O}_{o=1}$ identifying the region that the robot can use to interact with the object,
@@
$\mathcal{T}$ can be expressed through natural language [84, 85], an affordance category [13, 86, 87, 88], a point in 2D [89, 90], or another image of the object of interest [91, 92, 93]. With this formulation, affordance grounding and one-shot methods using prior information can be grouped together.
TABLE IV: Comparison of affordance grounding methods.
@@
We summarise the characteristics of affordance grounding methods in Table IV. Formulating the task as an additional input to the model enables affordance grounding methods to tackle generalization to object categories while avoiding an explicit object detection phase. Methods for explainability (Class Activation Maps [96]) highlight the region in the image that corresponds to the action [13, 86, 91, 95]. However, these regions are not bounded by object contours, limiting the application of these methods to unoccluded object settings. One-shot-based methods use an image as a prior to select objects of interest [97, 98], or segment affordance regions [92], based on the similarity between the input images and the prior (query image). However, the support image is assumed to be similar to the query images, thus implying that the object category in the scene should be known in advance.
@@
To cope with the limited amount of training images, methods adapt pre-trained models [84, 87, 88, 90], using knowledge transfer from video to image [86, 93] or from exocentric to egocentric views of the object [13, 91, 94]. In particular, multimodal models help generalising to unknown object categories or unknown actions (open vocabulary). For example, AffordanceCLIP adapts CLIP [35] with a learnable feature pyramid network to predict the affordance probability map [90]. A contrastive loss encourages the alignment between pixel-level embeddings within the annotated mask of the object and language features. AffordanceLLM [84] processes vision and language information to predict affordance segmentation tokens. The LLM generates text tokens encoding the object part used to perform the task and a mask token that is combined with the visual tokens using a transformer decoder to predict the affordance map. KBAG-Net [85] fuses language features extracted using BERT [99] with low and high resolution features from a visual backbone [33]. A convolutional decoder processes fused features to predict the affordance map.
Few of the methods [13, 91, 94] for affordance grounding focused on learning object affordances by building correspondences from the exocentric view of an object (human using the object) to the egocentric view (object only). Both LOCATE [91] and Cross-View-AG [13] during training combined a loss to learn the affordance category with losses to preserve the similarity between the feature maps of the exocentric and egocentric views. CALNet [94] models the correspondence between contact regions in the exocentric and egocentric views, concatenating the human keypoint features extracted from the exocentric view with the visual features of the egocentric perspective. Instead of learning directly from images, methods like Demo2Vec [93] and Hotspots [86] learn to transfer the affordance from videos of humans interacting with objects in household environments, e.g. oven, fridge, washing machine, to the images containing only the objects. Although these methods [13, 86, 91, 93, 94] can learn the object affordances from examples showing humans that perform actions, the egocentric views are composed by the object on a background without occlusions or clutter, limiting the generalisation to in-the-wild images.
@@
Despite generalising to different actions or task formulations, affordance grounding methods output confidence maps that are not bounded by object edges. The confidence maps could also overlap with other objects in case of clutter or with a human hand if the object is hand-held. Using a coarse confidence map when interacting with an object can lead an agent to misplace its hand, thus undermining the success of the interaction or harming the human.
III-D Hand pose estimation and synthesis
@@
To perceive the visual affordance, the agent predicts also how to perform the interaction with the object, i.e. the pose of the agent hand. Previous works [2, 8, 16, 17, 18, 100, 101, 102, 103] related the problem mostly to grasping rather than to visual affordances, and redefined the problem based on the hand: grasping detection for two-finger grippers [2, 16, 17, 18, 102], hand-object pose estimation for human hand [5, 8] (e.g. MANO model [24]), multi-finger grasping for three fingers Barrett hand [100]. Given the hand model, the image of the object, and the task, the model predicts the pose $P$ of the hand on the object,
@@
Predicting the pose of the agent’s hand, however, is challenging because the hand is not observed in the image, and therefore only the visual features of the object can be used.
Grasping detection. Assuming a two-finger gripper, the pose estimation is reformulated as prediction of grasping points directly on the image, encoding the parameters of the gripper as an oriented rectangle [2]. In particular, 1 DoF encodes the rotation with respect to the horizontal axis, 2 DoF encode the translation of the gripper centre (horizontal and vertical), and 2 DoF encode the geometry (opening width and fingers height). The underlying assumption is the availability of a depth map to obtain the full 7 DoF representation of the gripper in 3D (translation, rotation, and opening width). Given an RGB-D image $I\in\mathbb{R}^{W\times H\times 4}$, the model predicts a set of $G$ oriented rectangles $\{r_{g}\}^{G}_{g=1}$ with $r\in\mathbb{R}^{5}$ consisting of the rectangle centre coordinates, the rectangle size (width and height), and the orientation. Predicting the pose of a two-finger gripper on an object is challenging because each object has multiple grasping points, but only a subset of grasping poses leads to successful grasping. Moreover, when estimating the grasping points from a single view, only a side of the object is visible, limiting the number of feasible grasping points.
TABLE V: Comparison of grasping detection models.
@@
Table V summarises methods for grasping detection. Most of the methods [2, 16, 17, 18, 102] use RGB-D images to predict grasping rectangles, as the depth information provides geometric cues. Visual information is fused in different ways: in the first layers of the model [2, 16, 17, 102] (early fusion), or using a separate backbone to process RGB and depth before fusion (middle fusion) [18]. However, there are no results showing that a fusion mechanism is more effective than the others. The feature extraction is performed mainly by convolutional networks like ResNet [16, 17, 18, 103] or AlexNet [101] pre-trained on ImageNet [104], to transfer the features learned on large scale datasets. Methods can be categorised into single-stage and two-stage: single-stage methods [18, 101, 102] predict the final oriented rectangles from the image, either directly regressing the rectangle [18, 101] or considering the rectangle as a by-product of object segmentation [102]; two-stage methods [2, 16, 17, 103] first predict grasping candidates (coarse estimation) and then refine the predictions (fine estimation). The majority of two-stage methods adapt works for object detection (e.g. Faster R-CNN [47]) to grasping detection in different ways: separating the learning of the object location and grasp locations [17]; separating the learning of the quantized angle from the learning of the centre, width and height of the grasping rectangle [16]; separating the coarse prediction of grasping rectangles from the refinement based on the object segmentation [103]. Auxiliary tasks, such as object detection [16] and segmentation [103] constrain the prediction of the grasping rectangle to the object, reducing mistakes in cluttered scenes or when the object is not in foreground and completely visible. Other auxiliary tasks are: the likelihood of an image patch (non-overlapping piece of the image) containing a grasp [101] limiting the prediction of grasping rectangles to some parts of the image; and the grasping region segmentation [102] constraining the grasping rectangle to graspable region of the object, e.g. the handle of a spoon.
@@
Grasping detection formulation considers only the interaction of picking, resulting in non-functional solutions as the agent can grasp the object at any surface location. For example, the rim of a cup filled with liquid (suggesting the affordance of pouring the content) might be selected as a potential grasping point without considering that the liquid might be spilled or damage the robotic hand. Most of the methods for grasping detection [2, 16, 17, 18, 101, 105] assume that objects are observed on a tabletop or on the floor (top-down camera view). Hence, models fail to generalise to scenarios with different camera view-points or with occlusions.
@@
Multi-finger pose estimation and interaction synthesis. Previous works [5, 8, 100] considered as visual affordance the pose of an agent’s hand on objects in the scene. Given an image $I$, the model predicts the 6D pose of the hand on the object $\{[R|T]_{o}\}^{O}_{o=1}$ with $[R|T]$ representing pose of the hand and renders an image of the hand, $\tilde{I}\in\mathbb{R}^{W\times H\times 3}$ (interaction synthesis), showing how and where the hand interacts with the object, not what action is performed.
@@
TABLE VI: Comparison of multi-finger pose estimation and interaction synthesis methods.
@@
Table VI compares the characteristics of methods for multi-finger pose estimation and interaction synthesis. Methods are based on a coarse-to-fine approach locating first where the hand will interact with the object and then refining the pose using generative adversarial networks [8, 100] or diffusion models [5]. GanHand [8] estimates objects’ shapes and locations using an object 6D pose estimator or a reconstruction network. GanHand localises the object projecting its shape in the image plane and predicts the grasp type, i.e. the type of interaction between hand and object. The network predicts the coarse pose of the hand from the grasp type and the visual features, and refines the hand parameters to obtain the final shapes and poses (i.e. MANO model [106]), learned by minimising an adversarial loss with a discriminator. Multi-FinGAN [100] adapts GanHand architecture to perform the pose estimation of the Barrett end-effector on the object in the image. Contrary to GanHand, Multi-FinGAN uses the object reconstruction only to refine the coarse pose of the end-effector. As a consequence, the method underperforms if multiple objects are present in the scene. AffordanceDiffusion [5] is a cascade of two diffusion models to generate the image of the hand interacting with the object in the image. The diffusion process uses a prior (forearm mask) composed by a circle representing the hand and a rectangle representing the forearm. For every diffusion step, the first model predicts the denoised forearm mask from the features of the forearm mask obtained in the previous step, the object image, and the forearm mask projected on the object image. The second model combines the layout mask (prior) with the object image to synthesise the interaction.